OpenStudy (anonymous):

You throw a ball downward from a window at a speed of 2.0 m/s. How fast will it be moving when it hits the sidewalk 2.5 m below?

OpenStudy (anonymous):

this will require the use of kinematic equations: we have: Vstarting = 2.0 m/s HeightStarting = 2.5 m and we need to find Vfinal the kinematic equation of V^2 = V^2 + 2ah can be used velocity final ^2 = velocity starting ^2 + 2 * acceleration of gravity * change of height there are two terms which we are given, velocity starting and change in height, we are missing two terms: final velocity (which we are looking for) and acceleration of gravity. acceleration of gravity is a number which is always the same, and can be found usually in the front or back cover of a physics book. we are using meters, so we will use 9.81 m/sec^2 so now we have 3 knowns, and one unknown. all that is left is to plug the three known values into the kinematic equation I listed above, and use algebra to solve for the one unknown, which is final velocity.

OpenStudy (anonymous):

i guess it would be better to use s=ut+1/2at2

OpenStudy (anonymous):

and find t

OpenStudy (anonymous):

then use v=u+at

OpenStudy (anonymous):

where s is the distance=2.5 u=initial speed=0 v=final speed which we have to find a=acceleration which is due to gravity (take as positive) t=time

OpenStudy (anonymous):

if you want i will post the answer in 5 mins

OpenStudy (loser66):

 @dan815